In a Newton ring experiment the diameter of a eight dark ring is 0.72cm. If the the radius of curvature of plano-convex lense is 300cm. Calculate wave length of light
1
Expert's answer
2018-10-03T06:53:09-0400
For the dark ring system:
(d_n^2)/4R=nλ
The wave length of light is
λ=(d_n^2)/4nR=(0.72)^2/4(8)(300) =0.000054 cm=540 nm.